# On the Kantorovich contraction of Markov semigroups

### v1

@@ 1.2 Kantorovich semi-distances
The Kantorovich semi-distance is related to several well-known measures of distance between probability distributions. Specifically, the quantity $D_{\psi_{S}}(\mu_{1},\mu_{2})$ is the Kantorovitch distance between $\mu_{1}$ and $\mu_{2}$ (a.k.a. the Wasserstein 1-distance) while for $p\geq 1$ the Wasserstein $p$-distance $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})$ is defined by $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})=D_{\psi_{S}^{p}}(\mu_{1},\mu_{2})^{p}$. We also have the equality $D_{\varphi_{0}}(\mu_{1},\mu_{2})=\|\mu_{1}-\mu_{2}\|_{tv}$, that is $D_{\varphi_{0}}(\mu_{1},\mu_{2})$ is the total variation distance between the probability measures $\mu_{1}$ and $\mu_{2}$, and the weighted total variation norm $\|\mu_{1}-\mu_{2}\|_{V}$ is related to the Kantorovich semi-distance (1.6) by the formula
@@
$D_{\phi}(\mu_{1},\mu_{2})\leq(\mu_{1}\otimes\mu_{2})(\phi)\leq\mu_{1}(V)+\mu_{2}(V)<\infty$
@@ 2.1 Assumptions
$K(W)\leq\epsilon~V+c\quad\mbox{\rm and}\quad L(V)\leq\epsilon~W+c$ (2.1)
for some constants $\epsilon_{0}\in]0,1[$ and $c_{0}\geq 0$, then $P$ is a Markov $V$-operator and (H1) holds with $\epsilon=\epsilon_{0}^{2}$ and $c=(1+\epsilon_{0})c_{0}$.
@@ 2.6 Exponential decay of the $(\varphi_{V},\phi)$-Dobrushin contraction coefficient
with the semi-distance $\kappa_{\iota,V}$ as defined in (1.5). Moreover, if $\beta_{\kappa,\varphi_{0}}(P)<\infty$ then, for any $V$-semi-distance $\phi$, there exists a constant $c^{\prime}\geq 0$ such that
@@
It is also worth noting that the second part of Theorem 2.6 implies that if (H2) holds for the discrete metric (that is with $\kappa=\varphi_{0}$) then the $V$-Dobrushin contraction coefficient $\beta_{\varphi_{V}}(P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$.
@@ 2.7 Discussion
A contraction result for a continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ is obtained in [43], Theorem 4.8. Assuming the existence of a Lyapunov function $V$, this theorem shows that, for a symmetric and semi-distance $\kappa$ bounded by one, and with $\iota=1/2$ and $\rho=1$, the inequality $\beta_{\kappa_{\iota,\rho}}(Q_{t})<1$ holds for $t$ sufficiently large. However, this result further assumes that there exists a $t^{\prime}>0$ such that $\sup_{x\neq y}\ell_{\kappa}(Q_{t^{\prime}})(x,y)<1$ which, as proved in Lemma 1.1, is equivalent to assuming that $\beta_{\kappa}(Q_{t^{\prime}})<1$. In the context of $\mathbb{R}^{n}$-valued diffusions, an exponential contraction estimate is obtained in [30], assuming the existence of a Lyapunov function $V$ satisfying the geometric drift condition (2.27) and some additional key growth conditions (see Assumptions 2.3 in [30]). Specifically, Theorem 2.1 in [30] shows that $\beta_{\phi}(Q_{t})\leq\lambda^{t}$ for some constant $\lambda\in]0,1[$, where $\phi(x,y)=\varpi(\|x-y\|)+\varphi_{\rho}(x,y)$ for some judiciously chosen bounded non-decreasing concave continuous functions $\varpi$ and constant $\rho>0$. The proof of this result combines diffusion reflection coupling techniques with a general Ito-Tanaka type formula and martingales local times techniques to estimate the time spent by the processes on different level sets. By contrast, for the same diffusion model as the one considered in [30] and without any additional assumptions, exponential contraction estimates can be easily deduced from Theorem 2.6, following the discussion in Section 2.6.2 (see Section 3.5 for details). We however stress that these estimates do not establish that each Markov operator $Q_{t}$ is a contraction, unlike the result obtained in [30] (again, see Section 3.5 for details).
@@ 3.1 Some illustrative examples
3.1.1 Markov chains on the unit open interval with bounded densities
We let $S=]0,1[$, $\psi_{e}(x,y)=|x-y|$ be the Euclidean metric and, for any $\iota\in]0,1]$ we let
$\psi_{\iota}(x,y):=\left|\frac{1}{x}-\frac{1}{y}\right|^{\iota}+\left|\frac{1}{1-x}-\frac{1}{1-y}\right|^{\iota}.$ (3.3)
We note that $(S,\psi_{\iota})$ is a complete and separable metric space for any $\iota\in]0,1]$ while the space $(S,\psi_{e})$ is separable but not complete.
Next, given a probability measure $\nu(dy)=g(y)dy\in\mathcal{P}(S)$ having a bounded density $g$ and a Markov transition $Q(x,dy)=q(x,y)dy$ on ]0,1[]0,1[ such that $\|q\|<\infty$, we let
$\begin{array}[]{l}P(x,dy)=x\,Q(x,dy)+(1-x)\nu(dy)\end{array}$ (3.4)
and, for some $\iota\in]0,1[$, we define the continuous function
@@ 3.1.1 Markov chains on the unit open interval with bounded densities
Noting that $W_{\iota}\in{\cal B}_{\infty}(S)$ while $\|P(W_{\iota})\|\leq c:=2(\|q\|+\|g\|)/(1-\iota)<\infty$, it follows that (3.1) holds with $\Theta=1/W_{\iota}$ and thus, by (3.2), condition (H1) holds with $V=W_{\iota}$ (observe that $W_{\iota}$ is bounded below away from zero).
Assume first that the two probability distributions $\nu(dy)$ and $Q(x,dy)$ have disjoint supports for all $x\in S$. In this scenario, we have $\|\delta_{x}P-\delta_{0}P\|_{\tiny tv}=x$ implying that $\beta_{\varphi_{0}}(P)=1$ and therefore (H2) does not hold with the discrete metric, that is with $\kappa=\varphi_{0}$.
Assume now that, for all $x\in S$, the two probability distributions $\nu(dy)$ and $Q(x,dy)$ have $S$ as support, and assume also that their densities are continuous and strictly positive. In this scenario, (H2) holds with $\kappa=\varphi_{0}$ by Lemma 2.1, and Theorem 2.6 therefore ensures that the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ vanishes exponentially quickly as $n\rightarrow\infty$ for any $V$-semi-distance $\phi$. In particular, noting that the complete metric $\psi_{S}=\psi_{\iota}$ is a $V$-semi-distance function, Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
@@
Lastly, following an example proposed in [24], Section 2.1, we assume that
@@
so that the Markov transition $P$ has an unbounded density. In this context, we let $\iota\in]0,1/2[$ and show in Section 4.7 that there exist some parameters $\epsilon\in[0,1[$ and $c\geq 0$ such that
@@ 3.2 Markov chains on domains with a boundary
In this subsection we consider a Markov transition $P(x,dy)=p(x,y)dy$ with a uniformly bounded, strictly positive and continuous density $p(x,y)$ on a open and connected set $S\subset\mathbb{R}^{n}$ having a non-empty boundary $\partial S$. We define the continuous function
$d_{n}(x,\partial S):=\inf{\left\{\|x-y\|~:~y\in\partial S\right\}}$
and, for some $\iota\in]0,1]$, we equip $S$ with the metric
$\psi_{S}(x,y):=\left|\frac{1}{d_{n}(x,\partial S)}-\frac{1}{d_{n}(y,\partial S)}\right|^{\iota}+\|x-y\|^{\iota}.$ (3.10)
@@ 3.2.1 Bounded domains with a locally Lipschitz boundary
Assume that $S$ is bounded and has a locally Lipschitz boundary $\partial S$, and let $W(x)=d_{n}(x,\partial S)^{-\iota}+\|x\|$. In this scenario, we clearly have $W\in{\cal B}_{\infty}(S)$ and, since the function $W$ is integrable over the set $S$ (c.f. Lemma 5.2 in [4]), under the above assumptions on $P$ we have $\|P(W)\|<\infty$. Consequently, (3.1) is met with $\Theta=1/W$ and it follows from (3.2) that (H1) holds with $V=1+W$. On the other hand, since $P$ is assumed to have a strictly positive and continuous density, condition (H2) holds with $\kappa=\varphi_{0}$ by Lemma 2.1. By Theorem 2.6, it follows that for any $V$-semi-distance $\phi$ the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ converges to zero exponentially quickly as $n\rightarrow\infty$. In particular, one can check that the metric $\psi_{S}$ defined in (3.10) is the $V$-semi-distance function, and thus Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
@@ 3.2.2 Domains with a locally Lipschitz bounded boundary
Assume that $S$ is not necessarily bounded but has a locally Lipschitz bounded boundary $\partial S$. Assume also that there exists a function $W_{1}\in{\cal B}_{\infty}(\mathbb{R}^{n})$ such that $\sup_{x\in S}\|x\|^{\iota}/W_{1}(x)<\infty$ and such that $P(W_{1})\leq\epsilon_{1}W_{1}+c_{1}$ for some parameters $\epsilon_{1}\in[0,1[$ and $c_{1}>0$. In this scenario, condition (H1) holds with $V=1+W_{1}$ but, for this definition of $V$, the metric $\psi_{S}$ defined in (3.10) may not be a $V$-semi-distance. To solve this problem we let $W_{2}(x)=d_{n}(x,\partial S)^{-\iota}\in{\cal B}_{\infty}(S)$ and $c_{2}=\|P(W_{2})\|\leq 1+\|P(W_{2}1_{W_{2}\geq 1})\|$. Using the fact that $\{x\in S:W_{2}(x)\geq 1\}=\{x\in S:d_{n}(x,\partial S)\leq 1\}$ with $\partial S$ a bounded set, and recalling that the density $p(x,y)$ is assumed to be uniformly bounded, by using a similar argument as in the proof of Lemma 5.2 in [4] we can check that $\|P(W_{2}1_{W_{2}\geq 1})\|<\infty$. This implies that $c_{2}<\infty$ and thus, since
@@ 3.3.2 Lyapunov functions
Assume that $(g,h)$ and $(m_{g},m_{h})$ are continuous functions. In addition, assume that there exist some parameters $a_{1},b_{1}\in\mathbb{R}$, $a_{2},b_{2}>0$ and $q\geq p>0$ such that
@@ 3.4.2 A general approach to construct Lyapunov functions
$W(x):=\frac{1}{2}~\exp{(\delta\|x\|)}\quad\mbox{\rm and}\quad\Theta(x):=a_{1}~\exp{\left(-\delta(1-\lambda_{0})~\|x\|\right)}$
@@
Assume now that $\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)<\infty$ for some $p\geq 1$. In this scenario, letting $c_{\lambda,p}=c_{\lambda}+\sup_{x\in S}\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)^{1/p}$ we observe that
@@ 3.4.3 Example 1: Random iterated functions with local dissipative conditions
Consider the functions $(F,F_{z},G_{z})$ defined in Section 3.4.2 with ${\cal Z}=\mathbb{R}^{n}=S$ and
@@
Specifically, if $Z$ is a Gaussian random variable it follows from the calculations in Section 3.4.2 that (H1) is met when $V=(1/2)\exp(\delta\|x\|)$ for some $\delta>0$. If we further assume that the function $b$ is continuous, it follows from the discussion of Section 3.4.1 that condition (H2) is met with $\kappa=\varphi_{0}$. In this context, the calculations in Section 2.6.1 show that, for any $p\geq 1$ and with $\psi_{S}(x,y)=\|x-y\|$, the following exponential contraction estimate for the Wasserstein p-distance holds:

### v2

@@ 1.2 Kantorovich semi-distances
The Kantorovich semi-distance is related to several well-known measures of distance between probability distributions. Specifically, the quantity $D_{\psi_{S}}(\mu_{1},\mu_{2})$ is the Kantorovitch distance between $\mu_{1}$ and $\mu_{2}$ (a.k.a. the Wasserstein 1-distance) while for $p\geq 1$ the Wasserstein $p$-distance $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})$ is defined by $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})^{p}=D_{\psi_{S}^{p}}(\mu_{1},\mu_{2})$. We also have the equality $D_{\varphi_{0}}(\mu_{1},\mu_{2})=\|\mu_{1}-\mu_{2}\|_{tv}$, that is $D_{\varphi_{0}}(\mu_{1},\mu_{2})$ is the total variation distance between the probability measures $\mu_{1}$ and $\mu_{2}$, and the weighted total variation norm $\|\mu_{1}-\mu_{2}\|_{V}$ is related to the Kantorovich semi-distance (1.6) by the formula
@@
$D_{\phi}(\mu_{1},\mu_{2})\leq(\mu_{1}\otimes\mu_{2})(\phi)\leq\|\phi/\varpi_{V}\|\left(\mu_{1}(V)+\mu_{2}(V)\right)<\infty$
@@ 2.1 Assumptions
$K(W)\leq\epsilon_{1}~V+c_{1}\quad\mbox{\rm and}\quad L(V)\leq\epsilon_{2}~W+c_{2}$ (2.1)
for some constants $c_{1},c_{2}\geq 0$ and $\epsilon_{1},\epsilon_{2}\in]0,1[$, then $P$ is a Markov $V$-operator and (H1) holds with $\epsilon=\epsilon_{1}\epsilon_{2}$ and $c=(c_{1}\epsilon_{2}+c_{2})$.
@@ 2.6 Exponential decay of the $(\varphi_{V},\phi)$-Dobrushin contraction coefficient
with the semi-distance $\kappa_{\iota,V}$ as defined in (1.5). If, in addition, we have $\beta_{\kappa,\varphi_{0}}(P)<\infty$, then, for any $V$-semi-distance $\phi$, there exists a constant $c^{\prime}\geq 0$ such that
@@
It is also worth noting that the second part of Theorem 2.6 implies that, under (H1) and if (H2) holds for the discrete metric (that is with $\kappa=\varphi_{0}$), then, for any $V$-semi-distance $\phi$, the $(\varphi_{V},\phi)$-Dobrushin contraction coefficient $\beta_{\varphi_{V},\phi}(P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$.
@@ 2.7 Discussion
A contraction result for a continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ is obtained in [43], Theorem 4.8. Assuming the existence of a Lyapunov function $V$, this theorem shows that, for a symmetric and semi-distance $\kappa$ bounded by one, and with $\iota=1/2$ and $\rho=1$, the inequality $\beta_{\kappa_{\iota,\rho}}(Q_{t})<1$ holds for $t$ sufficiently large. However, this result further assumes that there exists a $t^{\prime}>0$ such that $\sup_{x\neq y}\ell_{\kappa}(Q_{t^{\prime}})(x,y)<1$ which, as proved in Lemma 1.1, is equivalent to assuming that $\beta_{\kappa}(Q_{t^{\prime}})<1$. In the context of $\mathbb{R}^{m}$-valued diffusions, an exponential contraction estimate is obtained in [30], assuming the existence of a Lyapunov function $V$ satisfying the geometric drift condition (2.27) and some additional key growth conditions (see Assumptions 2.3 in [30]). Specifically, Theorem 2.1 in [30] shows that $\beta_{\phi}(Q_{t})\leq\lambda^{t}$ for some constant $\lambda\in]0,1[$, where $\phi(x,y)=\varpi(\|x-y\|)+\varphi_{\rho}(x,y)$ for some judiciously chosen bounded non-decreasing concave continuous functions $\varpi$ and constant $\rho>0$. The proof of this result combines diffusion reflection coupling techniques with a general Ito-Tanaka type formula and martingales local times techniques to estimate the time spent by the processes on different level sets. By contrast, for the same diffusion model as the one considered in [30] and without any additional assumptions, exponential contraction estimates can be easily deduced from Theorem 2.6, following the discussion in Section 2.6.2 (see Section 3.5 for details). We however stress that these estimates do not establish that each Markov operator $Q_{t}$ is a contraction, unlike the result obtained in [30] (again, see Section 3.5 for details).
@@ 3.1 Some illustrative examples
3.1.1 Markov chains on the unit open interval
We let $S=]0,1[$, $\psi_{e}(x,y)=|x-y|$ be the Euclidean metric, $\iota\in]0,1[$,
$\psi_{\iota}(x,y):=\left|\frac{1}{x}-\frac{1}{y}\right|^{\iota}+\left|\frac{1}{1-x}-\frac{1}{1-y}\right|^{\iota}$ (3.3)
and we define the continuous function
@@ 3.1.1 Markov chains on the unit open interval
We note that $(S,\psi_{\iota})$ is a complete and separable metric space, while the space $(S,\psi_{e})$ is separable but not complete.
We now let $\nu(dy)=g(y)dy\in\mathcal{P}(S)$ and $\mu(dy)=h(y)dy\in\mathcal{P}(S)$ be probability measures with bounded densities, i.e. such that $\|g\|+\|h\|<\infty$, and we assume first that
$\begin{array}[]{l}P(x,dy)=x\,\mu(dy)+(1-x)\nu(dy).\end{array}$ (3.5)
Noting that $W_{\iota}\in{\cal B}_{\infty}(S)$ while $\|P(W_{\iota})\|\leq c:=2(\|g\|+\|h\|)/(1-\iota)<\infty$, it follows that (3.1) holds with $\Theta=1/W_{\iota}$ and thus, by (3.2), condition (H1) holds with $V=W_{\iota}$ (observe that $W_{\iota}$ is bounded below away from zero).
We further assume that the two probability distributions $\nu$ and $\mu$ have disjoint supports. In this situation, for any $x_{1},x_{2}\in S$ we have
$\|\delta_{x_{1}}P-\delta_{x_{2}}P\|_{\tiny tv}=|x_{1}-x_{2}|~\|\mu-\nu\|_{\tiny tv}=|x_{1}-x_{2}|$
implying that $\beta_{\varphi_{0}}(P)=1$. However, noting that,
$W_{\iota}\leq r\implies P(x,dy)\geq(1-\epsilon(r))\nu(dy),\quad\epsilon(r):=\sup_{W_{\iota}(x)\leq r}x\in]0,1[$
it follows that the local minorization condition (2.2) holds for any $r_{0}>0$ and with $\nu_{r}=\nu$. Hence, condition (H2) is met with the discrete metric, that is with $\kappa=\varphi_{0}$. In this context, by Theorem 2.6, for any $V$-semi-distance $\phi$ the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$. In particular, noting that the complete metric $\psi_{S}=\psi_{\iota}$ is a $V$-semi-distance function, Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
@@
Following an example proposed in [24], Section 2.1, we now assume that
@@
so that the Markov transition $P$ has an unbounded density. In this situation, we impose that $\iota\in]0,1/2[$ and show in Section 4.7 that there exist some parameters $\epsilon\in[0,1[$ and $c\geq 0$ such that
@@ 3.2 Markov chains on domains with a boundary
In this subsection we consider a Markov transition $P(x,dy)=p(x,y)dy$ with a uniformly bounded, strictly positive and continuous density $p(x,y)$ on a open and connected set $S\subset\mathbb{R}^{m}$ having a non-empty boundary $\partial S$. We define the continuous function
$d_{m}(x,\partial S):=\inf{\left\{\|x-y\|~:~y\in\partial S\right\}}$
and, for some $\iota\in]0,1[$, we equip $S$ with the metric
$\psi_{S}(x,y):=\left|\frac{1}{d_{m}(x,\partial S)}-\frac{1}{d_{m}(y,\partial S)}\right|^{\iota}+\|x-y\|^{\iota}.$ (3.10)
@@ 3.2.1 Bounded domains with a locally Lipschitz boundary
Assume that $S$ is bounded and has a locally Lipschitz boundary $\partial S$, and let $W(x)=d_{m}(x,\partial S)^{-\iota}+\|x\|^{\iota}$. In this scenario, we clearly have $W\in{\cal B}_{\infty}(S)$ and, since the function $W$ is integrable over the set $S$ (c.f. Lemma 5.2 in [4]), under the above assumptions on $P$ we have $\|P(W)\|<\infty$. Consequently, (3.1) is met with $\Theta=1/W$ and it follows from (3.2) that (H1) holds with $V=1+W$. On the other hand, since $P$ is assumed to have a strictly positive and continuous density, condition (H2) holds with $\kappa=\varphi_{0}$ by Lemma 2.1. By Theorem 2.6, it follows that for any $V$-semi-distance $\phi$ the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$. In particular, one can check that the metric $\psi_{S}$ defined in (3.10) is the $V$-semi-distance function, and thus Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
@@ 3.2.2 Domains with a locally Lipschitz bounded boundary
Assume that $S$ is not necessarily bounded but has a locally Lipschitz bounded boundary $\partial S$. Assume also that there exists a function $W_{1}\in{\cal B}_{\infty}(\mathbb{R}^{m})$ such that $\sup_{x\in S}\|x\|^{\iota}/W_{1}(x)<\infty$ and such that $P(W_{1})\leq\epsilon_{1}W_{1}+c_{1}$ for some parameters $\epsilon_{1}\in[0,1[$ and $c_{1}>0$. In this scenario, condition (H1) holds with $V=1+W_{1}$ but, for this definition of $V$, the metric $\psi_{S}$ defined in (3.10) may not be a $V$-semi-distance. To solve this problem we let $W_{2}(x)=d_{m}(x,\partial S)^{-\iota}$ and $c_{2}=\|P(W_{2})\|\leq 1+\|P(W_{2}1_{W_{2}\geq 1})\|$. Using the fact that $\{x\in S:W_{2}(x)\geq 1\}=\{x\in S:d_{m}(x,\partial S)\leq 1\}$ with $\partial S$ a bounded set, and recalling that the density $p(x,y)$ is assumed to be uniformly bounded, by using a similar argument as in the proof of Lemma 5.2 in [4] we can check that $\|P(W_{2}1_{W_{2}\geq 1})\|<\infty$. This implies that $c_{2}<\infty$ and thus, since
@@ 3.3.2 Lyapunov functions
Assume that $(g,h)$ and $(m_{g},m_{h})$ are continuous functions on $S=\mathbb{R}^{d}$. In addition, assume that there exist some parameters $a_{1},b_{1}\in\mathbb{R}$, $a_{2},b_{2}>0$ and $q\geq p>0$ such that
@@ 3.4.2 A general approach to construct Lyapunov functions
$W(x):=\exp{(\delta\|x\|)}\quad\mbox{\rm and}\quad\Theta(x):=a_{1}~\exp{\left(-\delta(1-\lambda_{0})~\|x\|\right)}$
@@
Assume now that $\sup_{x\in S}\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)<\infty$ for some $p\geq 1$. In this scenario, letting $c_{\lambda,p}=c_{\lambda}+\sup_{x\in S}\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)^{1/p}$ we observe that
@@ 3.4.3 Example 1: Random iterated functions with local dissipative conditions
Consider the functions $(F,F_{z},G_{z})$ defined in Section 3.4.2 with ${\cal Z}=\mathbb{R}^{m}=S$ and
@@
Specifically, if $Z$ is a Gaussian random variable it follows from the calculations in Section 3.4.2 that (H1) is met when $V=\exp(\delta\|x\|)$ for some $\delta>0$. If we further assume that the function $b$ is continuous, it follows from the discussion of Section 3.4.1 that condition (H2) is met with $\kappa=\varphi_{0}$. In this context, the calculations in Section 2.6.1 show that, for any $p\geq 1$ and with $\psi_{S}(x,y)=\|x-y\|$, the following exponential contraction estimate for the Wasserstein p-distance holds: